OpenStudy (anonymous):

if you have an unknown amount of both 1% and 4% milk. what is the ratio to make 5 gallons of 2% milk?

OpenStudy (dumbcow):

they have 4% milk?? anyway let x,y be amount of 1%,4% milk respectively x+y = 5 .01x + .04y = .02*5=.1 trick to get rid of decimals, multiply both equations by 100 100x + 100y = 500 x + 4y = 10 Solve with substitution x = 10-4y 100(10-4y) +100y = 500 1000 - 400y +100y =500 1000 -300y = 500 -300y = -500 y = 5/3 x = 10-4*(5/3) = 10-20/3 = 10/3 so mixture is made with 10/3 gallons 1% and 5/3 gallons 4%